increase in value at a rate of 5% compounded annually for the next 5 years. How much will the coin be worth at the end of the fifth year? 1. Today Meghan Brown purchased an investment grade gold coin for $200,000. She expects it to I/Y PV PMT FV

Today Meghan Brown purchased an investment grade gold coin for $200,000. She expects it to increase in value at a rate of 5% compounded annually for the next 5 years. How much will the coin be worth at the end of the fifth year? 
Answer #1

These are financial calculator functions

put values as follow

N=5

I/Y=5

PV=200000

PMT=0

Click CPT

Click FV=255256.31 is the answer..

using another method as

FV=PV*(1+r)^n

=200000*(1+5%)^5

=255256.31

the above is answer..